# Distributional Random Forests for Complex Survey Designs on Reproducing Kernel Hilbert Spaces

### v1

@@ Abstract
We study estimation of the conditional law $P(Y|X=\mathbf{x})$ and continuous functionals $\Psi(P(Y|X=\mathbf{x}))$ when $Y\in\mathcal{Y}$ takes values in a locally compact Polish space, $X\in\mathbb{R}^{p}$, and the observations arise from a complex survey design. We propose a survey-calibrated distributional random forest (SDRF) that incorporates complex-design features via a pseudo-population bootstrap, PSU-level honesty, and a Maximum Mean Discrepancy (MMD) split criterion computed from kernel mean embeddings of Hájek-type (design-weighted) node distributions. We provide a framework for analyzing forest-style estimators under survey designs; establish design consistency for the finite-population target and model consistency for the super-population target under explicit conditions on the design, kernel, resampling multipliers, and tree partitions. As far as we are aware, these are the first results on model-free estimation of conditional distributions under survey designs. Simulations under a stratified two-stage cluster design provide finite sample performance and demonstrate the statistical error price of ignoring the survey design. The broad applicability of SDRF is demonstrated using NHANES: We estimate the tolerance regions of the conditional joint distribution of two diabetes biomarkers, illustrating how distributional heterogeneity can support subgroup-specific risk profiling for diabetes mellitus in the U.S. population.
@@ 1 Introduction and Motivation
Experimental design is a key driver of successful statistical practice [97]. A rigorous study design protects and makes transparent the assumptions that link the data to scientific conclusions, thus improving the reliability and reproducibility of the empirical findings [71].
Complex survey design [42] methods provide a powerful methodology for population-representative inference in a wide range of settings, such as biomedical studies [59]. In the finite-population regime [25], a probability sample with known inclusion probabilities and carefully constructed weights allows design-unbiased (or asymptotically design-unbiased) estimation of population functionals. These estimators can exhibit satisfactory statistical properties, such as efficiency, provided that appropriate statistical calibration and inference procedures are used. This design-based framework is particularly compelling when paired with randomized interventions: probability surveys deliver external validity (representativeness of a well-defined target population), while randomized experiments deliver internal validity (causal identification under the design). Together, they support the policy-relevant generalization of causal effects [45].
Large, nationally representative health studies, such as the National Health and Nutrition Examination Survey (NHANES), which monitors health and nutrition in the U.S. population, are canonical examples of the impact of complex survey methodology in practice [68]. When design information is properly incorporated and analyzed with appropriate statistical methods, such data can be used to quantify disease burden, monitor disparities, and inform clinical and public health policy in a cost-effective way, all while ensuring valid population inference.
@@
Despite the central role of complex survey designs in nationally representative studies, predictive tools for complex surveys remain limited in two main respects. First, most existing methods target simple functionals such as conditional means or medians, rather than the full conditional distribution, and are typically based on linear models or classical nonparametric smoothers such as Nadaraya–Watson estimators and spline-based methods [61]. Second, available procedures often presuppose low-dimensional covariates and focus on smoothing over relatively simple or very smooth function classes. By contrast, the use of modern machine-learning methods that can capture complex structures more effectively in high-dimensional settings such as neural networks [5, 98] or tree-based ensembles [26]-—within the survey context remains relatively scarce [66, 67].
@@
From an application’s perspective, many scientific questions demand distributional targets beyond means or medians [50], including multivariate quantiles, tolerance regions [34], tail probabilities, and other functionals of multivariate conditional laws. In NHANES, for example, multivariate distribution estimates can be used to define tolerance regions that detect outlying biomarker profiles given subject-specific characteristics. Treating complex-survey observations as i.i.d. data discards design information, degrading calibration and invalidating design-based uncertainty quantification, while properly incorporating weights and design features preserves both consistency and representativeness.
@@
Modern machine-learning methods, such as neural networks and regression-tree ensembles, are effective in many high-dimensional regimes and can capture intricate statistical associations that are difficult to model with classical nonparametric techniques. However, they are typically formulated under i.i.d. sampling and, when naively applied to complex surveys, can distort node-wise conditional distributions, bias risk estimates, and compromise inference. A design-aware approach must propagate weights and clustering through splitting and aggregation, respect replicate-weight or bootstrap schemes that mimic the original selection, and produce distributional outputs for vector-valued $Y$, not merely conditional means. These requirements call for distributional learning procedures that integrate survey design into both the objective function and the algorithmic construction.
@@
From a machine-learning perspective, integrating predictive algorithms with complex survey designs creates opportunities to increase the statistical efficiency of estimators, mitigate selection bias, reduce overfitting through design-guided regularization, improve subgroup fairness, and achieve proper predictive calibration in inference tasks in applied settings.
@@
Our paper. We study nonparametric estimation of multivariate conditional distributions under complex survey designs. Let $(Y,X)$ denote a pair consisting of a vector-valued response $Y\in\mathcal{Y}$ and covariates $X\in\mathbb{R}^{p}$ observed through a probability sample drawn under a complex survey design. We introduce a design-aware distributional random forest estimator $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ that combines survey-weighted tree construction and aggregation with loss functions induced by mean kernel embeddings of probability measures in a separable reproducing kernel Hilbert space. This yields a single forest-based procedure that (i) targets the full conditional law $\mathbb{P}(Y\mid X=\mathbf{x})$ and, as plug-in estimators, a broad class of component-wise and joint functionals such as conditional quantiles, cross-moments, covariance matrices and tail probabilities, and (ii) remains consistent under the joint law induced by the super–population model and the survey design.
@@
On the theoretical side, we formulate the estimation problem in terms of the maximum mean discrepancy (MMD) [37, 80, 89], a family of statistical distances that metrizes weak convergence in suitable classes of probability measures, and show that the forest weighing the survey achieves design and model consistency; that is, the MMD between $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ and $\mathbb{P}_{Y\mid X=\mathbf{x}}$ converges uniformly to zero over $x$ at rates governed by tree-partition and design complexity. We obtain a decomposition of the MMD risk into design, algorithmic, and stochastic terms, prove that each component vanishes under mild moment conditions on inclusion and replicate weights. Our analysis quantifies the impact of the number of trees $B$ and the bootstrap or replicate-weight draws $M$ on node-level predictions and on the variance of the ensemble, showing a reduction in variance of order $O(M^{-1/2})$ while preserving the consistency of the design. We emphasize that, even in the i.i.d. setting, the theory of classical random forests is far from trivial [9, 10] and has required substantial effort to develop since Breiman’s original proposal of random forests [17]. Moreover, many questions remain open, for instance regarding finite-sample guarantees [55] or extensions to dependent data [36].
@@
Our work is based on two strands. On the survey side, foundational contributions make the sampling mechanism explicit through survey weights [96, 84, 40, 77, 30, 82, 81, 39], including the Horvitz–Thompson and Hájek estimators [44, 28], as well as design-aware procedures widely used, for instance, in model checking, hypothesis testing, and regression [51, 75, 61, 74, 70, 47]. On the machine-learning side, we exploit the statistical learning properties of random forests [17] to reduce variance estimation and even bias in many regimes via the bagging step [15, 55]. Random forests [17] and their distributional variants (through kernel mean embeddings and MMD) provide flexible nonparametric tools that achieve competitive performance in high-dimensional settings across many benchmarks [31] and can be used efficiently for two-sample testing, variable selection, and clustering, in addition to predictive tasks. However, existing random forest algorithms are largely developed for i.i.d. data   design-aware random forest methodology and theory for complex surveys remain underdeveloped.
@@
From a regression-model perspective, existing survey methods focus primarily on parametric and semiparametric estimation using (pseudo) probability weights and estimating equations related to M- and Z-estimation [91, 69]. Classical nonparametric smoothers—including local polynomials [48], Nadaraya–Watson estimators [43], and spline-based additive models—have been adapted to complex surveys [14]. However, most of this literature focuses on continuous outcomes [83, 46] and conditional learning; comparatively less attention has been paid to multivariate outcomes and full conditional learning.
@@
From a theoretical perspective, design-based central limit theorems are available for a variety of survey designs [6, 60, 78, 11]. Empirical-process formulations replace the uniform empirical measure with Horvitz–Thompson/Hájek-weighted measures; under suitable weight conditions, the resulting processes inherit many i.i.d. properties. Survey-aware resampling is well–developed [76]: bootstrap and replicate-weight methods with asymptotic guaranties approximate sampling distributions and can be adapted to distributional functionals. However, existing theoretical developments are not directly tailored to distributional random forests under complex designs, where the interaction between algorithmic randomness, survey weights, and distributional loss functions requires a novel theoretical survey framework. In this paper, we provide a design-aware distributional random forest for survey data, together with a corresponding theoretical framework for its design-based and model-based properties. When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential for statistical consistency, as we illustrate in Table S3.
@@ 2 Set-up and Background
Suppose that we observe a sample $D=\{(X_{i},Y_{i},w_{i})\}_{i=1}^{n_{s}}$ drawn from a finite population $D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ of size $N$ under a sampling design $\mathbf{p}$ that may depend on auxiliary design variables $Z_{i}$, which we will formalize shortly. Here, $Y_{i}\in\mathcal{Y}$ is the outcome, $X_{i}\in\mathbb{R}^{p}$ are covariates and $w_{i}\in\mathbb{R}^{+}$ are survey weights (possibly incorporating nonresponse adjustment and calibration). Let $P^{N}_{Y\mid X=\mathbf{x}}$ denote the conditional distribution of a finite population induced by $D^{N}$. Let $P_{Y\mid X=\mathbf{x}}$ denote the super-population conditional law. Consider a fixed covariate value $\mathbf{x}\in\mathbb{R}^{p}$ and a measurable functional $\Psi:\mathcal{M}_{b}(\mathcal{Y})\to(\mathcal{V},d_{v})$ taking some probability measures on $\mathcal{Y}$ in some separable metric space $(\mathcal{V},d_{v})$, and it is continuous at the target. Our objective is to construct an estimator $\widehat{\Psi}_{N}(\mathbf{x})$ that is design-consistent for the finite-population target $\Psi(P^{N}_{Y\mid X=\mathbf{x}})$ and model-consistent for the super-population target $\Psi(P_{Y\mid X=\mathbf{x}})$. We will work on the joint probability space $(\Omega\times\mathcal{S}^{N},\,\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\Omega\times\mathcal{S}^{N}})$. Specifically, denote $P_{\mathcal{S}^{N}|\omega}$ the measure in the design space, we want that for any $\varepsilon>0$,
$P_{S^{N}\mid\omega}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P^{N}_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{$P_{\Omega}$--a.e.\,(design consistency), and }$
@@ 2.1 Survey Design
We follow the finite population framework considered by [13, 41, 42]: Consider a sequence of finite populations $\{\mathcal{U}^{N}\}_{N\geq 1}$. Let $U^{N}:=\{1,\dots,N\}$, the index set for each $\mathcal{U}^{N}$ and $\mathcal{S}^{N}:=\{S:S\subset U^{N}\}$ the collection of subsets of $U^{N}$. Let $\sigma(\mathcal{S}^{N})$ denote the $\sigma$-algebra generated by $\mathcal{S}^{N}$. We view the finite population $\mathcal{U}^{N}$ as a realization of i.i.d. random vectors $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ in a common probability space $(\Omega,\mathcal{A},P)$, where $Z_{i}\in\mathcal{Z}$ denotes the auxiliary (design) information available in the frame. Let $D^{N}:=\{(X_{i},Y_{i})\}_{i=1}^{N}$ denote the finite-population data and write $Z^{N}:=(Z_{1},\dots,Z_{N})$. For any region ${A}(\mathbf{x})\subset\mathcal{X}$ indexed by some fixed $\mathbf{x}\in\mathcal{X}$, denote the empirical conditional measure induced by ${D}^{N}$ as
@@
${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}I(X_{i}\in{A}(\mathbf{x}))\;\delta_{Y_{i}},\text{ where $\delta_{Y_{i}}$ the Dirac measure,}$
@@
$N(\mathbf{x})=\sum_{i\in[N]}I(X_{i}\in A(\mathbf{x}))$. A sampling scheme can be formalized as a transition function $\bm{p}:\Omega\times\sigma(\mathcal{S}^{N})\mapsto[0,1]$. From $\bm{p}$, fix any $\omega\in\Omega$, we define a probability measure in the design space $(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}))$ as $P_{\mathcal{S}^{N}}(A):=\sum_{S\in A}\bm{p}({Z}^{N}(\omega),S)$, $\forall A\in\sigma(\mathcal{S}^{N})$. We can then uniquely define a product probability measure in the product space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}))$ as
@@
$P_{\Omega\times\mathcal{S}^{N}}(E\times S):=\int_{E}\bm{p}(\omega,\{S\}){P}(d\omega),\hskip 8.5359pt\forall E\times\{S\}\text{ in simple rectangles in }\mathcal{A}\otimes\sigma(\mathcal{S}^{N}).$
@@
Thus, a random sample $S\sim P_{\mathcal{S}^{N}}$ can be obtained by ${\xi}^{N}=(\xi_{1},\dots,\xi_{N})$, a random vector on $\mathcal{S}^{N}$, where $\xi_{i}:=\mathbbm{1}_{(i\in S)}\in\{0,1\}$. Define first-order inclusion probabilities $\pi_{i}(\omega):=\mathbb{E}_{\mathcal{S}}(\xi_{i}|Z(\omega))=\sum_{\{S:i\in S\}}\bm{p}(S,{Z}^{N}(\omega))$, and second-order inclusion probabilities $\pi_{ij}(\omega):=\mathbb{E}(\xi_{i}\xi_{j}|Z(\omega))$. Denote $n_{s}:=card(S)=\sum_{i=1}^{N}\xi_{i}$, the size of the random sample $S$ that is a random variable. We do not have access to the entire finite population of size $N$, but we work with a sample of size $n_{s}$. This leads to estimators that use weights to construct consistent estimators ([56]). For example, the Hájek (HJ) estimator for the true conditional probability ${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})$ in the design space is
$\widehat{P}^{N}_{HJ,Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{\widehat{N}(\mathbf{x})}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x}))\;\delta_{Y_{i}},\quad\widehat{N}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x})).$
@@
The SDRF estimator, which we denote as $\widehat{\Psi}_{N}(\mathbf{x})$, adopts the form Hájek, but with critical modifications from design-compatible resampling and an adaptive data-driven partition $L^{*}(\mathbf{x})$ in place of $A(\mathbf{x})$. We assume the following with respect to the design.
@@
    Stable sampling fraction: $n_{s}/N\rightarrow_{p_{d}}f$ for some $f\in(0,1]$, and $P_{\mathcal{S}^{N}|\omega}(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$, with some $\kappa>4$.
@@
Assumptions (D1)–(D4), or minor variants thereof, are standard regularity conditions in the survey-sampling literature. (D2)–(D4) imply the design-based law of large numbers and central limit theorem statements, summarized in Lemma 2.1. These probabilistic bounds will be invoked repeatedly to establish the consistency of $\widehat{\Psi}_{N}(\mathbf{x})$. (D1) allows us to separate the super–population randomness from the design randomness and to treat the design expectations conditionally on $(X^{N},Z^{N})$. (D2) imposes a stable sampling fraction and rules out designs with a decreasing sample size. (D3) assumes that the inclusion probabilities of first order are uniformly of order $n/N$ and bounded away from 0 and $1$. This prevents extreme survey weights. (D4) is common [12, 11, 19] and a mild second order condition that controls the dependence among the inclusion indicators $\{\xi_{i}\}_{i=1}^{N}$ to be of order $1/n_{s}$.

### v2

@@ Abstract
We study estimation of the conditional law $P(Y|X=\mathbf{x})$ and continuous functionals $\Psi(P(Y|X=\mathbf{x}))$ when $Y\in\mathcal{Y}$ takes values in a locally compact Polish space, $X\in\mathbb{R}^{p}$, and the observations arise from a complex survey design. We propose a survey-calibrated distributional random forest (SDRF) that incorporates complex-design features via a pseudo-population bootstrap, PSU-level honesty, and a Maximum Mean Discrepancy (MMD) split criterion computed from kernel mean embeddings of Hájek-type (design-weighted) node distributions. We provide a framework for analyzing forest-style estimators under various survey designs; establish design consistency for the finite-population target and model consistency for the super-population target under explicit conditions on the design, kernel, resampling multipliers, and tree partitions. As far as we are aware, these are the first results on model-free estimation of conditional distributions under survey designs. Simulations under a stratified two-stage cluster design provide finite sample performance and demonstrate the statistical error price of ignoring the survey design. The broad applicability of SDRF is demonstrated using NHANES: We estimate the tolerance regions of the conditional joint distribution of two diabetes biomarkers, illustrating how distributional heterogeneity can support subgroup-specific risk profiling for diabetes mellitus in the U.S. population.
@@ 1 Introduction and Motivation
Experimental design is a key driver of successful statistical practice yates1964sir. A rigorous study design protects and makes transparent the assumptions that link the data to scientific conclusions, thus improving the reliability and reproducibility of the empirical findings peng2015reproducibility.
Complex survey design Han_Weller_2021 methods provide a powerful methodology for population-representative inference in a wide range of settings, such as biomedical studies lumley2011complex. In the finite-population regime DeMatteis2025, a probability sample with known inclusion probabilities and carefully constructed weights allows design-unbiased (or asymptotically design-unbiased) estimation of population functionals. These estimators can exhibit satisfactory statistical properties, such as efficiency, provided that appropriate statistical calibration and inference procedures are used. This design-based framework is particularly compelling when paired with randomized interventions: probability surveys deliver external validity (representativeness of a well-defined target population), while randomized experiments deliver internal validity (causal identification under the design). Together, they support the policy-relevant generalization of causal effects cinelli2025challengesstatisticsdozenchallenges; imbens2015causal.
Large, nationally representative health studies, such as the National Health and Nutrition Examination Survey (NHANES), which monitors health and nutrition in the U.S. population, are canonical examples of the impact of complex survey methodology in practice mozumdar2011persistent. When design information is properly incorporated and analyzed with appropriate statistical methods, such data can be used to quantify disease burden, monitor disparities, and inform clinical and public health policy in a cost-effective way, all while ensuring valid population inference.
@@
Despite the central role of complex survey designs in nationally representative studies, predictive tools for complex surveys remain limited in two main respects. First, most existing methods target simple functionals such as conditional means or medians, rather than the full conditional distribution, and are typically based on linear models or classical nonparametric smoothers such as Nadaraya–Watson estimators and spline-based methods lumley2017fitting. Second, available procedures often presuppose low-dimensional covariates and focus on smoothing over relatively simple or very smooth function classes. By contrast, modern machine-learning methods that can capture complex structure in high-dimensional settings—such as neural networks BauerKohler2019DeepLearningCurseDimensionality; zammit2022deep and tree-based ensembles diaz2006gene—remain relatively uncommon in survey applications matabuena2025screeningdiabetesmellitusus; matabuena2025kernel.
@@
In scientific applications, many questions demand distributional targets beyond means or medians kneib2023rage, including multivariate quantiles, tolerance regions fraser1953nonparametric, tail probabilities, and other functionals of multivariate conditional laws. In NHANES, for example, multivariate distribution estimates can be used to define tolerance regions that detect outlying biomarker profiles given subject-specific characteristics. Treating complex-survey observations as i.i.d. data discards design information, degrading calibration and invalidating design-based uncertainty quantification, while properly incorporating weights and design features preserves both consistency and representativeness.
@@
Modern machine-learning methods, such as neural networks and regression-tree ensembles, are effective in many high-dimensional regimes and can capture intricate statistical associations that are difficult to model with classical nonparametric techniques. However, they are typically formulated under i.i.d. sampling assumptions and, when naively applied to complex surveys, can distort node-wise conditional distributions, bias risk estimates, and compromise inference. A design-aware approach must propagate weights and clustering through splitting and aggregation, respect replicate-weight or bootstrap schemes that mimic the original selection, and produce distributional outputs for vector-valued $Y$, not merely conditional means. These requirements call for distributional learning procedures that integrate survey design into both the objective function and the algorithmic construction.
@@
From a machine-learning perspective, integrating predictive algorithms with complex survey designs creates opportunities to increase the statistical efficiency of estimators, mitigate selection bias, reduce overfitting through design-guided regularization, improve subgroup fairness, and achieve proper predictive calibration in inference tasks in applied settings
@@
Our paper: Overview. We study nonparametric estimation of multivariate conditional distributions under complex survey designs. Let $(Y,X)$ denote a pair consisting of a vector-valued response $Y\in\mathcal{Y}$ and covariates $X\in\mathcal{X}\subseteq\mathbb{R}^{p}$ observed through a probability sample drawn under a complex survey design. We introduce a design-aware distributional random forest estimator $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ that combines survey-weighted tree construction and aggregation with loss functions induced by mean kernel embeddings of probability measures in a separable reproducing kernel Hilbert space. This yields a single forest-based procedure that (i) targets the full conditional law $\mathbb{P}(Y\mid X=\mathbf{x})$ and, as plug-in estimators, a broad class of component-wise and joint functionals such as conditional quantiles, cross-moments, covariance matrices and tail probabilities, and (ii) remains consistent under the joint law induced by the super–population model and the survey design. By “full conditional law” we mean the entire probability distribution $\mathbb{P}(Y\mid X=\mathbf{x})$, not only summary functionals such as conditional means $\mathbb{E}[Y\mid X=\mathbf{x}]$ or quantiles. When $Y$ is multivariate, we estimate the joint conditional distribution, preserving dependence structure among components.
@@
On the theoretical side, we formulate the estimation problem in terms of the maximum mean discrepancy (MMD) gretton2012kernel; sejdinovic_equivalence_2013; szekely2003statistics, a family of statistical distances that metrizes weak convergence in suitable classes of probability measures, and show that the proposed design-aware forest achieves design and model consistency; that is, the MMD between $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ and $\mathbb{P}_{Y\mid X=\mathbf{x}}$ converges uniformly to zero over $x$ at rates governed by tree-partition and design complexity. We obtain a decomposition of the MMD risk into design, algorithmic, and stochastic terms, prove that each component vanishes under mild moment conditions on inclusion and replicate weights. Crucially, we show that PSU-level sample splitting combined with pseudopopulation bootstrap preserves both honesty and design-based consistency, despite the dependency structure induced by the complex sampling mechanism. We emphasize that, even in the i.i.d. setting, the theory of classical random forests is far from trivial biau2008consistency; biau2016random and has required substantial effort to develop since Breiman’s original proposal of random forests breiman2001random. Moreover, many questions remain open, for instance regarding finite-sample guarantees JMLR:v26:24-0255 or extensions to longitudinal data goehry2020random@book.
@@
Our work is based on two strands. On the survey side, foundational contributions make the sampling mechanism explicit through survey weights yates1953sampling; hajek1960limiting; roy1960estimating; ericson1969subjective; shindell1963population; hajek1959optimal, including the Horvitz–Thompson and Hájek estimators horvitz1952generalization; dupavc1975jaroslav, as well as design-aware procedures widely used, for instance, in model checking, hypothesis testing, and regression koch1975strategies; rao1987simple; lumley2017fitting; rabe2006multilevel; natarajan2012extension; jimenez2018estimation. On the machine-learning side, we exploit the statistical learning properties of random forests breiman2001random to reduce variance estimation and even bias in many regimes via the bagging step breiman1996bagging; JMLR:v26:24-0255. Random forests breiman2001random and their distributional variants (through kernel mean embeddings and MMD) provide flexible nonparametric tools that achieve competitive performance in high-dimensional settings across many benchmarks fernandez2014we and can be used efficiently for two-sample testing, variable selection, and clustering, in addition to predictive tasks. However, existing random forest algorithms are largely developed for i.i.d. data   design-aware random forest methodology and theory for complex surveys remain underdeveloped.
@@
From a regression-model perspective, existing survey methods focus primarily on parametric and semiparametric estimation using (pseudo) probability weights and estimating equations related to M- and Z-estimation van1996m; nan2013general. Classical nonparametric smoothers—including local polynomials kikechi2017local, Nadaraya–Watson estimators harms2010kernel, and spline-based additive models—have been adapted to complex surveys breidt2005model. However, most of this literature focuses on continuous outcomes signorino2018using; iparragirre2023variable and conditional learning; comparatively less attention has been paid to multivariate outcomes and full conditional learning.
@@
From a theoretical perspective, design-based central limit theorems are available for a variety of survey designs bellhouse2001central; lumley2012empirical; schochet2022design; boistard2017functional. Empirical-process formulations replace the uniform empirical measure with Horvitz–Thompson/Hájek-weighted measures; under suitable weight conditions, the resulting processes inherit many i.i.d. properties. Survey-aware resampling is well–developed rao1988resampling: bootstrap and replicate-weight methods with asymptotic guaranties approximate sampling distributions and can be adapted to distributional functionals. However, existing theoretical developments are not directly tailored to distributional random forests under complex designs, where the interaction between algorithmic randomness, survey weights, and distributional loss functions requires a novel theoretical survey framework. In this paper, we provide a design-aware distributional random forest for survey data, together with a corresponding theoretical framework for its design-based and model-based properties. When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential for statistical consistency, as we illustrate in Section S2 of the Supplementary Material.
@@ 2 Set-up and Background
For notational convenience, for any $a\in\mathbb{Z}_{>0}$, we will denote $a]:=\{1,\dots,a\}$. ‘:=’ will be used to emphasize definition. Suppose that we observe a sample $D=\{(X_{i},Y_{i},w_{i})\}_{i\in[n_{s}]}$ drawn from a finite population $D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i\in[N]}$ of size $N$ under a sampling design $\mathbf{p}$ that may depend on auxiliary design variables $Z_{i}$, which we will formalize shortly. Here, $Y_{i}\in\mathcal{Y}$ is the outcome, $X_{i}\in\mathbb{R}^{p}$ are covariates and $w_{i}\in\mathbb{R}^{+}$ are survey weights (possibly incorporating nonresponse adjustment and calibration). Let $P^{N}_{Y\mid X=\mathbf{x}}$ denote the conditional distribution of a finite population induced by $D^{N}$. Let $P_{Y\mid X=\mathbf{x}}$ denote the super-population conditional law. Consider a fixed covariate value $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ and a measurable functional $\Psi:\mathcal{M}_{b}(\mathcal{Y})\to(\mathcal{V},d_{v})$ taking on some probability measures on $\mathcal{Y}$ in some separable metric space $(\mathcal{V},d_{v})$, and which is continuous at the target. Our objective is to construct an estimator $\widehat{\Psi}_{N}(\mathbf{x})$ that is design-consistent for the finite-population target $\Psi(P^{N}_{Y\mid X=\mathbf{x}})$ and model-consistent for the super-population target $\Psi(P_{Y\mid X=\mathbf{x}})$. We will work on the joint probability space $(\Omega\times\mathcal{S}^{N},\,\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\Omega\times\mathcal{S}^{N}})$. Specifically, denoting $P_{\mathcal{S}^{N}|\omega}$ as the measure in the design space, we want that for any $\varepsilon>0$,
$P_{\mathcal{S}^{N}\mid\omega}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P^{N}_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{$P_{\Omega}$--a.e.\,(design consistency), and }$
@@ 2.1 Survey Design
We follow the finite population framework considered by boistard_functional_2017; Han_Weller_2019; Han_Weller_2021: Consider a sequence of finite populations $\{\mathcal{U}^{N}\}_{N\geq 1}$. Let $U^{N}:=\{1,\dots,N\}$, the index set for each $\mathcal{U}^{N}$ and $\mathcal{S}^{N}:=\{S:S\subset U^{N}\}$ the collection of subsets of $U^{N}$. Let $\sigma(\mathcal{S}^{N})$ denote the $\sigma$-algebra generated by $\mathcal{S}^{N}$. We view the finite population $\mathcal{U}^{N}$ as a realization of i.i.d. random vectors $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ in a common probability space $(\Omega,\mathcal{A},P)$, where $Z_{i}\in\mathcal{Z}$ denotes the auxiliary (design) information available in the frame. Let $D^{N}:=\{(X_{i},Y_{i})\}_{i=1}^{N}$ denote the finite-population data and write $Z^{N}:=(Z_{1},\dots,Z_{N})$. For any region ${A}(\mathbf{x})\subset\mathcal{X}$ indexed by some fixed $\mathbf{x}\in\mathcal{X}$, denote the empirical conditional measure induced by ${D}^{N}$ as
@@
${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in{A}(\mathbf{x}))\;\delta_{Y_{i}},\text{ where $\delta_{Y_{i}}$ the Dirac measure,}$
@@
$N(\mathbf{x})=\sum_{i\in[N]}\mathbbm{1}(X_{i}\in A(\mathbf{x}))$. A sampling scheme can be formalized as a transition function $\mathbf{p}:\Omega\times\sigma(\mathcal{S}^{N})\mapsto[0,1]$. Given $\mathbf{p}$, fix any $\omega\in\Omega$ and define a probability measure in the design space $(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}))$ as $P_{\mathcal{S}^{N}}(A):=\sum_{S\in A}\mathbf{p}({Z}^{N}(\omega),S)$, $\forall A\in\sigma(\mathcal{S}^{N})$. We can then uniquely define a product probability measure in the product space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}))$ as
@@
$P_{\Omega\times\mathcal{S}^{N}}(E\times S):=\int_{E}\mathbf{p}(\omega,\{S\}){P}(d\omega),\hskip 8.5359pt\forall E\times\{S\}\text{ in simple rectangles in }\mathcal{A}\otimes\sigma(\mathcal{S}^{N}).$
@@
Thus, a random sample $S\sim P_{\mathcal{S}^{N}|\omega}$ can be obtained by ${\xi}^{N}=(\xi_{1},\dots,\xi_{N})$, a random vector on $\mathcal{S}^{N}$, where $\xi_{i}:=\mathbbm{1}_{(i\in S)}\in\{0,1\}$. Define first-order inclusion probabilities $\pi_{i}(\omega):=\mathbb{E}_{\mathcal{S}}(\xi_{i}|Z(\omega))=\sum_{\{S:i\in S\}}\mathbf{p}(S,{Z}^{N}(\omega))$, and second-order inclusion probabilities $\pi_{ij}(\omega):=\mathbb{E}(\xi_{i}\xi_{j}|Z(\omega))$. Denote $n_{s}:=card(S)=\sum_{i=1}^{N}\xi_{i}$ as the size of the random sample $S$. We do not have access to the entire finite population of size $N$, but we work with a sample of size $n_{s}$. This leads to estimators that use weights to construct consistent estimators (Lohr2021-cm). For example, the Hájek (HJ) estimator for the true conditional probability ${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})$ in the design space is
$\widehat{P}^{N}_{HJ,Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{\widehat{N}(\mathbf{x})}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in A(\mathbf{x}))\;\delta_{Y_{i}},\quad\widehat{N}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in A(\mathbf{x})).$
@@
The SDRF estimator, which we denote as $\widehat{\Psi}_{N}(\mathbf{x})$, adopts a Hájek-like form, but with critical modifications from design-compatible resampling and an adaptive data-driven partition $L^{*}(\mathbf{x})$ in place of $A(\mathbf{x})$. We assume the following with respect to the design.
@@
    Stable sampling fraction: $n_{s}/N\rightarrow_{p_{d}}f$ for some $f\in(0,1]$, and $\exists\delta\in(0,\;f)$, $C<\infty$, such that for $N$ large enough, $P_{\mathcal{S}^{N}|\omega}(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$, for some $\kappa>4$.
@@
Assumptions (D1)–(D4), or minor variants thereof, are standard regularity conditions in the survey-sampling literature. (D2)–(D4) imply the design-based law of large numbers and central limit theorem statements, summarized in Lemma 2.1. These probabilistic bounds will be invoked repeatedly to establish the consistency of $\widehat{\Psi}_{N}(\mathbf{x})$. (D1) allows us to separate the super–population randomness from the design randomness and to treat the design expectations conditionally on $(X^{N},Z^{N})$. (D2) imposes a stable sampling fraction and rules out designs with a decreasing sample size. (D3) assumes that the inclusion probabilities of first order are uniformly of order $n/N$ and bounded away from 0 and $1$. This prevents extreme survey weights. (D4) is common Boistard2012-gc; boistard2017functional; Cardot2010-cn and a mild second order condition that controls the dependence among the inclusion indicators $\{\xi_{i}\}_{i=1}^{N}$ to be of order $1/n_{s}$.